OpenStudy (anonymous):

Admission to the movies costs $8 for adults and $4 for kids. The total bill for your entire extended family was $72. This scenario can be modeled by the equation 8a + 4k = 72. Which is not a possible combination of adults and kids? 3 adults and 12 kids 5 adults and 6 kids 7 adults and 4 kids 8 adults and 2 kids please help?

OpenStudy (anonymous):

i think that it would be C

OpenStudy (anonymous):

because 7 times 8 and 4 times 4 addes up to 72

OpenStudy (johnweldon1993):

So you would replace the number of adults and kids in the choices into the equation...I'll do an example....for choice A 8a + 4k = 72 Choice 'a' says 3 adults and 12 kids So we replace every 'a' we see with 3 and every 'k' we see with 12...so 8(3) + 4(12) = 72 24 + 48 = 72 72 = 72 This IS a possible combination so it is NOT the answer

OpenStudy (johnweldon1993):

Remember to read the question...which is NOT a possible combination...so C cannot be correct because like you said....56 + 16 = 72...it works...so it is NOT the answer

OpenStudy (anonymous):

ughhhh.... ok let me redo this

OpenStudy (johnweldon1993):

Well from the 2 examples we already did...we know it isn't A or C lol...let me know what you get :)

OpenStudy (anonymous):

i got B

OpenStudy (johnweldon1993):

That would be correct! 40 + 24 does not equal 72 So good work! :)

OpenStudy (anonymous):

thank you :) could you help with this type of problem: Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (johnweldon1993):

This is a point slope form... \[\large y - y_0 = m(x - x_0)\] You replace 'm' with your slope replace y_o with your y value and replace x_o with your x value... so it is...?

OpenStudy (anonymous):

ummm

OpenStudy (johnweldon1993):

So we take that equation \[\large y - y_0 = m(x - x_0)\] you have a slope (-6) for which you replace 'm' with You have an x and a y value (-2 , -3) (x , y) so replace those too So we put those in the equation...and get....?

OpenStudy (anonymous):

okk! i think i understand how those ones are done!

OpenStudy (johnweldon1993):

What did you get for the answer choice?